OpenStudy (anonymous):

Help need HELP WILL FAN AND GVE MDL @cool12900 @linn99123 @31356 @Luigi0210 @whpalmer4 @Oscarmenot @toughguy

OpenStudy (anonymous):

@Bookworm14 @SkiTTleoooo47 @BellaNina865

OpenStudy (anonymous):

help with wat

OpenStudy (anonymous):

math so much lol

OpenStudy (anonymous):

@Snoopy_Lascano

OpenStudy (anonymous):

whats up im killer in math

OpenStudy (anonymous):

k can u hlp?

OpenStudy (anonymous):

The volume of 10,000 drops of a liquid is 10 fluid ounces. What is the volume of 100 drops?

OpenStudy (anonymous):

give me a sec ok

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

you got a gmail

OpenStudy (anonymous):

yes and a yahoo

OpenStudy (anonymous):

this is my yahoo lol

OpenStudy (anonymous):

whats your gmail address

OpenStudy (anonymous):

asomedude12@gmail.com

OpenStudy (anonymous):

your keegan

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

100=0.1 fl oz

OpenStudy (anonymous):

get on google

OpenStudy (anonymous):

thanks can u do mor? k

OpenStudy (anonymous):

ya but get on google

OpenStudy (anonymous):

k

OpenStudy (anonymous):

on google lol

OpenStudy (anonymous):

I suck in math, so I can't help you with any math questions

OpenStudy (anonymous):

do you see a mesage that shows a fire logo on it

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok send me a messag at zacwam@gmail.com

OpenStudy (anonymous):

k

OpenStudy (anonymous):

and also what do you need help with

OpenStudy (anonymous):

this The weight of 10 identical samples of a substance is 0.1 pound. What is the weight of 1000 samples?

OpenStudy (anonymous):

1000x = 10 so x = 10/1000 or 1/100 so 10 of these would be 10(1/100) = 1/10 = 0.1

OpenStudy (anonymous):

hey added you to my contact lol

OpenStudy (anonymous):

ok